OpenStudy (anonymous):

How Do I Add Fractions With Different Denominators? Ex: 1/8 + 1/6

OpenStudy (anonymous):

\[\frac{1}{8} * \frac {3}{3} +\frac{1}{6}*\frac{4}{4}\]

OpenStudy (anonymous):

you need to get common denominators first find what number both 8 and 6 go into, 8, 16 ,24 32..... and 6, 12,18,24. Then multiply the fractions both tops and bottoms

OpenStudy (anonymous):

8 gets multiplied by 3 because 24 is the 3rd multiple of 8 in the list and 6 gets multiplied by 4 for the same reason

OpenStudy (anonymous):

\[\frac{3}{24}+\frac{4}{24} =\frac{7}{24}\]

OpenStudy (anonymous):

when adding/subtracting fractions never change the denominator (bottom number)

OpenStudy (anonymous):

1/8 + 1/6 = (6*1+8*1) /(6*8) ==> 14/48 = 7/24